OpenStudy (anonymous):

Find two 4 digit natural numbers that multiply to 4 to the power of 8 + 6 to the power of 8 + 9 to the power of 8

OpenStudy (anonymous):

@kidspence did you see answer to last problem?

OpenStudy (anonymous):

Yeah I'm confused though

OpenStudy (anonymous):

what step?

OpenStudy (anonymous):

i cant understand it properly

OpenStudy (anonymous):

All.. :$ LOL

OpenStudy (anonymous):

Find two 4 digit natural numbers that multiply to equal 4 to the power of 8 + 6 to the power of 8 + 9 to the power of 8

OpenStudy (anonymous):

oh this one i don't know. i meant last one

OpenStudy (anonymous):

I know the last one I'm confused on

OpenStudy (anonymous):

Find the smallest positive whole number that when added to: 2001 x 2002 x 2004 x 2005 - 2003 x 2003 you get a square number

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

that one yes?

OpenStudy (anonymous):

ok let me start again. to make life easier put \[2003=x\]

OpenStudy (anonymous):

so \[2001=x-2\] \[2002=x-1\] \[2004=x+1\] \[2005=x+2\] clear so far?

OpenStudy (anonymous):

YUp

OpenStudy (anonymous):

now some algebra. we get \[(x-2)(x-1)(x+1)(x+2)-x^2\]

OpenStudy (anonymous):

\[(x-2)(x+2)(x-1)(x+1)-x^2\] \[(x^2-4)(x^2-1)-x^2\] \[x^4-5x^2+4-x^2\] \[x^4-6x^2+4\]

OpenStudy (anonymous):

that is what you have if x was 2003

OpenStudy (anonymous):

and this is not a perfect square, but if you add 5 you get \[x^4-6x^2+9\] and that is a perfect square. it is \[(x^2-3)^2\]

OpenStudy (anonymous):

so if you add 5 to that big mess above you get \[(2003^2-3)^2\] a perfect square

OpenStudy (anonymous):

we can check with a calculator and see that this is right, but the algebra does all the work for you

OpenStudy (anonymous):

yup, they are equal look here where i typed them both in together http://www.wolframalpha.com/input/?i=%282003^2-3%29^2%2C2001*2002*2004*2005-2003^2%2B5